# Instruction-Based Molecular Graph Generation with Unified Text-Graph Diffusion Model

### v1

@@ Abstract
Recent advancements in computational chemistry have increasingly focused on synthesizing molecules based on textual instructions. Integrating graph generation with these instructions is complex, leading most current methods to use molecular sequences with pre-trained large language models. In response to this challenge, we propose a novel framework, named UTGDiff (Unified Text-Graph Diffusion Model), which utilizes language models for discrete graph diffusion to generate molecular graphs from instructions. UTGDiff features a unified text-graph transformer as the denoising network, derived from pre-trained language models and minimally modified to process graph data through attention bias. Our experimental results demonstrate that UTGDiff consistently outperforms sequence-based baselines in tasks involving instruction-based molecule generation and editing, achieving superior performance with fewer parameters given an equivalent level of pretraining corpus. Our code is availble at https://github.com/ran1812/UTGDiff.
Introduction
@@
Molecules, as fundamental units of matter, possess structures and properties that determine the characteristics of substances. Research into novel molecules is pivotal for chemistry, biology, and drug discovery (Schneider and Fechner 2005; Hughes et al. 2011), as these fields depend on the discovery of molecules with specific properties. A significant challenge in drug discovery is designing molecules that exhibit desired features, maintain structural stability, and interact with specific targets (Hartenfeller and Schneider 2011). Traditionally, this process has been resource-intensive and time-consuming (Dickson and Gagnon 2009; Khanna 2012). However, recent advancements in deep learning have revolutionized drug design by offering cost-effective methods (Nag et al. 2022; Askr et al. 2023), drawing significant attention to the research of generating drug-like molecules (Bagal et al. 2021; Schneuing et al. 2022).
In this work, we aim to bridge the gap between molecular structures and natural language, designing a generative model that functions akin to a drug design expert (illustrated in Figure 1). By receiving specific instructions, this model can precisely generate candidate molecules with the indicated characteristics. We focus on tasks including instruction-based molecule generation (Edwards et al. 2022) and molecule editing (Fang et al. 2023), striving to integrate two modalities of information in drug design: aligning text descriptions to ensure desired properties and generating graph structures that represent potential molecules.
@@
Figure 1: A deep generative model that works like a drug design specialist, with instruction input and molecule output.
One widely adopted approach involves generating molecular strings from instructions, such as SMILES or SELFIES (Weininger 1988; Krenn et al. 2020), which serve as a stopgap for describing molecular graphs using language. Several recent studies have explored this method by employing auto-regressive language models  (Edwards et al. 2022; Liu et al. 2023d; Pei et al. 2023), deploying large language models  (Achiam et al. 2023; Ye et al. 2023; AI4Science and Quantum 2023), or modifying existing molecules based on specified instructions  (Liang et al. 2023).
However, despite these advancements, current methods face significant challenges. The primary limitation of molecular strings lies in their insufficient capability to accurately represent complex molecular structures, as it is widely recognized that they lack spatial structure information such as neighborhood (Jiang et al. 2022; Wu et al. 2023; Song et al. 2020), which could limit performance. Additionally, aligning the language space with the molecular space often results in deviations from natural molecular distributions, such as generating unstable or rare structures through simple string reassembly (Druchok et al. 2021). It is therefore essential to develop a multimodal model that possesses robust molecular structure representational capabilities while maintaining instruction-following abilities.
As a non-sequential data structure, graph generation capacity could be constrained by fixed orders imposed by auto-regressive models. This limitation motivates our usage of diffusion, a powerful non-sequential generative model. Therefore, we propose the Unified Text-Graph Diffusion Model (UTGDiff), a novel diffusion framework for text-guided graph generation. UTGDiff extends vanilla transformer as graph denoising networks through attention bias from graph edge categories, ensuring a strong capacity for learning graph structure representations and executing instruction texts without additional graph encoding modules. The equivariance of our model in graph structure ensures structural modeling abilities beyond asymmetric SMILES (O’Boyle and Dalke 2018; Arús-Pous et al. 2019).
@@
We performed extensive experiments on the pre-trained UTGDiff model across diverse tasks. Our results show that UTGDiff achieves improved outcomes over language model baselines in drug design tasks (instruction-based molecule generation and molecule editing) despite its small parameter size (125M). Additionally, our use of a diffusion-generation pipeline and text-graph denoising network achieves a relatively high accuracy of over 85% in terms of chemical validity. In summary, our contributions are:
  • We present a unified text-graph diffusion framework for instruction-based graph generation, enabling the creation of molecular structures that are instructed by text and represent accurate chemical structures.
  • We propose UTGDiff, which advances the state of the art by employing this unified text-graph transformer that effectively captures the features of both graph and text data. The model distinguishes itself from prior work by enhancing molecular representation and generation.
  • Experimental results demonstrate that UTGDiff outperforms existing language model baselines in several critical areas of drug design, including instruction-based molecule generation and molecule editing, with an equivalent level of pretraining corpus.
Related Works
@@
Early attempts at molecule generation introduced sequence-based models for SMILES strings (Dai et al. 2018; Gómez-Bombarelli et al. 2018). Due to the lack of structural information, recent studies have explored graph-based methods, including auto-regressive models (Grisoni et al. 2020; You et al. 2018), VAEs (Simonovsky and Komodakis 2018; Jin, Barzilay, and Jaakkola 2018; Liu et al. 2018), GANs (De Cao and Kipf 2018), and Normalizing Flows (Ma and Zhang 2021; Zang and Wang 2020). A relatively novel approach, offering superior performance and better scalability on larger molecules, is Graph Diffusion (Niu et al. 2020; Jo, Lee, and Hwang 2022; Vignac et al. 2022), which serves as our foundational framework.
@@
Conditional molecule generation, an important topic in molecule generation, aims to generate a molecular graph that satisfies a specific property. Early attempts are based on a scalar property such as dipole moment $\mu$ and p-logP scores (the octanol-water partition coefficients) (Li, Zhang, and Liu 2018; Huang et al. 2023). More complex conditions involve textual instructions, with existing solutions including sequence-based generation (Edwards et al. 2022; Christofidellis et al. 2023; Liu et al. 2023d; Li et al. 2023; Pei et al. 2024) and multimodal generation employing a text encoder and graph generative network (Su et al. 2022; Liu et al. 2024).
@@
In addition to the instruction-based cross-modal molecule generation models discussed above, some studies represent molecules and text in cross-modal models for molecule property prediction or text generation tasks (Liu et al. 2023c; Zhao et al. 2024; Liu et al. 2023a), employing contrastive alignment, cross-modal projector, or unified text-graph backbone.
@@
Previous Gaussian noise in continuous diffusion models has been shown to be insufficient for discrete data structure. As a solution, researchers introduced the discrete diffusion method, initially discussed in text diffusion (Austin et al. 2021; He et al. 2022) and later adapted for graph data (Vignac et al. 2022; Kong et al. 2023).
Preliminaries
Figure 2: Overview of the UTGDiff framework. It incorporates attention bias into the vanilla transformer, forming a unified text-graph transformer that serves as a denoising network for discrete graph diffusion to generate molecular graphs from instructions. Noise decays some nodes and edges into [MASK] during training (forward process), with the reverse process aiming to recover original graphs as the training objective. Sampling starts with a masked graph and iterates $T$ times to reduce noise.
The objective of the instruction-based molecule generation task is to generate a molecular graph $G=(V,E)$ from a given textual description $S$ specifying molecular functions and properties. The graph consists of $m$ nodes $V\in\mathbb{R}^{m\times a}$, which represent node attributes, and edges $E\in\mathbb{R}^{m\times m\times b}$, which represent edge attributes. Each attribute $v_{i}\in\mathbb{R}^{a}$ and $e_{i,j}\in\mathbb{R}^{b}$ is a one-hot vector representing the categories of nodes and edges, respectively. The instruction text is defined as $S=[s_{1},\dots,s_{n}]$, where $s_{i}$ denotes the $i$-th token, and $n$ is the length of the instruction. The goal is to perform conditional graph generation $V,E\sim P(V,E\mid S)$.
Method
Overview of Diffusion Framework
We propose a text-conditioned diffusion framework for molecular graph generation, as depicted in Figure 2. As a generative model, the diffusion model iteratively refines random noise into structured data through a series of transformations. It involves the forward process, which gradually adds noise to data via a noise model, and the reverse process, which denoises the data to recover the original structure using a denoising network. At each reverse step, the model estimates and removes the noise component, progressively generating the final structure. The core component of our framework is the novel unified text-graph denoising network, described in the Reverse Process section.
Forward Process and Noise Model
The forward process, associated with the noise model $q$, progressively adds noise to corrupt the clean graph $G^{0}$, creating an increasingly noisy graph sequence $(G^{1},\dots,G^{T})$:
$q(G^{1:T}\mid G^{0})$ $=\prod^{T}_{t=1}q(G^{t}\mid G^{t-1})$ (1)
Since traditional Gaussian noise is not suitable for graphs due to its disruption of sparsity and graph theory properties such as connectivity, we follow the discrete diffusion setting in (Vignac et al. 2022). The noise model $q$ is represented by a sequence of transition matrices $(Q^{1},\dots,Q^{T})$. It acts on a one-hot encoding $x\in\mathbb{R}^{d}$ over $d$ categories, where $[Q^{t}]_{i,j}$ indicates the probability of transferring the category $i$ to $j$ at time step $t$. We independently diffuse each one-hot category-based attribution in nodes $V$ and edges $E$ using separate matrices $Q_{V}^{t}$ and $Q_{E}^{t}$. The forward process is:
@@
It can also be expressed as:
$q(G^{t}\mid G^{0})=(V\bar{Q}_{V}^{t},E\bar{Q}_{E}^{t})$ (3)
$\bar{Q}_{V}^{t}=Q_{V}^{1}\dots Q_{V}^{t},\quad\bar{Q}_{E}^{t}=Q_{E}^{1}\dots Q_{E}^{t}$
@@
Specifically, we employ absorbing discrete diffusion (Austin et al. 2021), where each element $v_{i}^{t}$ (and $e_{i,j}^{t}$) independently decays into an absorbing state based on the noise schedule probabilities. The absorbing state is defined as a [MASK] token, shared by text and graph tokens. Following the noise schedule $\beta(t)=(T-t+1)^{-1}$, the corresponding transition matrix $Q^{t}$ is formally expressed as follows, applying to both nodes and edges simultaneously, where $z$ is the index for [MASK] token:
i,j =$\left\{\begin{array}[]{l l l}{{1}}&{{\mathrm{if}}}&{{i=j=z}}\\ {{1-\beta(t)}}&{{\mathrm{if}}}&{{i=j\not=z}}\\ {{\beta(t)}}&{{\mathrm{if}}}&{{j=z,i\not=z}}\end{array}\right.$ (4)
Reverse Process and Denoising Network
Overview
@@
The reverse process relies on a denoising neural network $\phi_{\theta}$ parameterized by $\theta$. This network is trained to invert the forward process by predicting $G^{t-1}$ from $G^{t}$ using the text instruction $S$, formally represented as:
$p_{\theta}(G^{0:T-1}\mid G^{T},S)=\prod^{T}_{t=1}p_{\theta}(G^{t-1}\mid G^{t},S).$ (5)
As the formula indicates, the denoising network must simultaneously encode the textual instruction and graph to generate the probability distributions of nodes and edges at next time step. To ensure sufficient comprehension of instructions, we retain the transformer model and propose the Unified Text-Graph Transformer as our denoising network.
Tokenizer
To adapt the vanilla transformer for processing graph data, we first address the representation of graph nodes by treating them as new tokens. The initial embeddings are denoted as $h^{0}=\text{Emb}(S+V)=[h_{1},\dots,h_{n},h_{n+1},\dots,h_{n+m}]\in\mathbb{R}^{(n+m)\times d_{h}}$, corresponding to $m$ graph nodes and $n$ text tokens, where $h_{1},\dots,h_{n}$ are encodings derived from the token list $S=[s_{1},\dots,s_{n}]$ and $h_{n+1},\dots,h_{n+m}$ are encodings based on the categorical list of atoms in the molecule (e.g., [[S],[O],[O]] for sulfur oxide). Our model treats each graph category as a novel token to prevent interference from irrelevant tokens, thereby avoiding misrepresentations such as interpreting cobalt as a combination of carbon and oxygen.
Attention Bias
Next, we address the representation of edges. In the attention layer, which includes parameters for values, keys, queries, and outputs, i.e., $W_{V},W_{Q},W_{K}\in\mathbb{R}^{d_{h}\times d_{k}},W_{O}\in\mathbb{R}^{d_{k}\times d_{h}}$, we introduce an additional bias term to incorporate edge information. The attention score between the $i$-th and $j$-th graph tokens is now formalized as:
$\hat{A}_{i,j}^{l}$ $=\frac{1}{\sqrt{d_{k}}}\left(h_{i}^{l}W_{Q}\right)\left(h_{j}^{l}W_{K}\right)^{{}^{\prime}}+b_{i,j}^{l},$ (6)
$A^{l}$ $=\mathrm{softmax}(\hat{A}^{l}),\ \text{Attn}(H)=A^{l}H^{l}W_{V}W_{O}$
Here, $b_{i,j}$ represents the bias added to the graph tokens to incorporate structural information (Ying et al. 2021; Choromanski et al. 2022; Park et al. 2022; Zhao et al. 2023). The most fundamental requirement for encoding graph data is the ability to distinguish whether two nodes in the graph are neighbors, which corresponds to the expressive power of a single-layer GNN (Xu et al. 2018). In addition, edge attribution must maintain symmetry for structural interpretability. Representing the initial embedding from the edge category as $b_{i,j}^{0}$, we therefore define the attention bias as follows:
$b_{i,j}^{l}=\begin{cases}b_{i,j}^{0},&\text{if }l=0\\ A_{i,j}^{l-1},&\text{otherwise}\\ \end{cases}$ (7)
In addition to fulfilling the structural requirements outlined above and theoretically discussed in Appendix B, a vital advantage is that our model requires minimal modifications to existing language models, enabling them to process graph and text data simultaneously without additional graph neural network and benefiting from initial pre-trained weights. Moreover, our model enhances the integration of text and graph data, as the layer-level interactions between modalities facilitate a better alignment, allowing text information to be incorporated into graph data at multiple scales.
Output
After obtaining the final hidden states $h^{L}$ for tokens and $b^{L}$ for edges through stacked attention layers, we map these hidden states back to the vocabulary space to determine the probability of each category for diffusion. We employ two separate masked language model heads for nodes and edges at the end of the transformer to decode these hidden states to the distribution of a clean graph.
Ultimately, we obtain the nodes logits $\log p_{\theta}(v_{i}\mid G^{t},S)$ and edges logits $\log p_{\theta}(e_{i,j}\mid G^{t},S)$ for a clean graph as output. Assembling these logits into matrices as $\hat{p}(V)\in\mathbb{R}^{m\times a}$, $\hat{p}(E)\in\mathbb{R}^{m\times m\times b}$ , we summarize our unified text-graph denoising network as:
$\phi_{\theta}(G^{t},S)=\hat{p}(G)=[\hat{p}(V),\hat{p}(E)]$ (8)
Sampling
There remains a gap between the distribution of the clean graph and the required next-step distribution as formula (5). This gap is fixed through the $x_{0}$-parameterization, which adds noise back to the clean graph prediction, represented as $p_{\theta}(G^{t-1}\mid G^{t},S)=q(G^{t-1}\mid G^{t},\phi_{\theta}(G^{t},S))$, yielding a distribution from which $G^{t-1}$ is sampled:
$p_{\theta}(v^{t-1}_{i}\mid G^{t},S)=\sum_{v\in\mathcal{V}}q(v^{t-1}_{i}\mid v_{i}=v,v^{t}_{i})\;{\hat{p}}_{i}(v)$ (9)
$p_{\theta}(e^{t-1}_{i,j}\mid G^{t},S)=\sum_{e\in\mathcal{E}}q(e^{t-1}_{i,j}\mid e_{i,j}=e,e^{t}_{i,j})\;{\hat{p}}_{i,j}(e)$
Now sampling can be performed as Algorithm 2. It begins with graph $G^{T}$, where each node and edge is masked. At each step $t$, it calculates the distribution and samples the graph at step $t-1$, gradually reducing the noise and eventually resulting in a clean graph $G^{0}$. Our model can implicitly sample the molecular length through the empty token, which is crucial as target molecule may be unique for instructions.
Training Objective
Since the denoising model takes a noisy graph $G^{t}$ as input and aims to predict the clean graph $G^{0}$, we optimize the cross-entropy loss (represented as $\mathrm{CE}$) for each node and edge between the predicted probabilities $\hat{p}(G)$ and the clean graph $G$ during training. The diffusion loss is defined as:
loss $=l(\hat{p}(V),V)+l(\hat{p}(E),E)$ (10)
$=\sum_{1\leq i\leq m}\mathrm{CE}(v_{i},\hat{p}_{i}(V))+\sum_{1\leq i,j\leq m}\mathrm{CE}(e_{i,j},\hat{p}_{i,j}(E))$
In addition, as previously discussed, maintaining balanced training across different modalities is crucial to prevent the model from diminishing its text comprehension ability after extensive training, as the tokens of the graph and text are non-overlapping. To address this issue, we incorporate a textual loss component to preserve the model’s text comprehension capability, which is defined as the cross-entropy loss for each token between the predicted logits $\hat{p}(S)$ and the ground-truth token in the instruction. Through the denoising model, the logits can be directly obtained alongside $\hat{p}(V)$ following a masked language model head, expressed as $\phi_{\theta}(G^{t},S)=[\hat{p}(S),\hat{p}(V),\hat{p}(E)]$. The final loss is thus represented as:
$\text{loss}=l(\hat{p}(V),V)+l(\hat{p}(E),E)+l(\hat{p}(S),S)$ (11)
See Algorithm 1 for the complete training procedure.
Algorithm 1 Training Algorithm
0:  A graph $G=(V,E)$ with description $S$
1:  repeat
2:     Sample $t\sim U(1,\dots,T)$
3:     Sample $G^{t}\sim V\bar{Q}_{V}^{t}\times E\bar{Q}_{E}^{t}$
4:     $\hat{p}(S),\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
5:     loss = $l_{CE}(\hat{p}(E),E)+l_{CE}(\hat{p}(V),V)+l_{CE}(\hat{p}(S),S)$
6:     optimizer.step(loss)
7:  until converged
Algorithm 2 Sampling Algorithm
1:  $G^{T}$ with all masked nodes and edges
2:  for $t=T$ to $1$ do
3:     $\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
7:  end for
8:  return  $G^{0}$
@@
Specialist Auto-regressive (w.o. pretrain) T5-small 77M 0.704 0.578 0.525 2.89 0.064 0.608
T5-base 248M 0.731 0.605 0.545 2.48 0.069 0.660
@@
Specialist Auto-regressive (pretrain) MolT5-small 77M 0.703 0.568 0.517 2.49 0.079 0.721
MolT5-base 248M 0.721 0.588 0.529 2.18 0.081 0.772
@@
LLM chatgpt3.5 (0-shot) 7B 0.703 0.568 0.517 2.49 0.079 0.721
Chatgpt3.5 (10-shot) 7B 0.847 0.708 0.624 0.57 0.139 0.887

### v2

@@ Abstract
Recent advancements in computational chemistry have increasingly emphasized generating and editing molecules from textual instructions. However, integrating graph generation with instruction understanding remains challenging, as most existing approaches either rely on molecular sequences in text modality with limited structural information, or struggle with multimodal alignment in graph diffusion methods. To address these limitations, we propose UTGDiff (Unified Text-Graph Diffusion Model), a novel framework that utilizes pre-trained language models for discrete graph diffusion, enabling the generation of molecular graphs from instructions. UTGDiff introduces a unified text-graph transformer as a denoising network, adapted with minimal modifications from language models to process graph data via attention bias. Experimental results show that UTGDiff consistently outperforms both sequence-based and conditional graph-diffusion baselines on instruction-based molecule generation and editing tasks with fewer parameters, covering instructions specifying molecular structures or properties.
[…]
4522
1 Introduction
@@
Molecules possess structures and properties that determine the characteristics of substances. Research into novel molecules is crucial for fields such as chemistry, biology, and drug discovery [43]. A central challenge in drug discovery is designing molecules that are stable, exhibit desired functionalities, and interact with specific biological targets [17]. Traditionally, this process has been resource-intensive and time-consuming [8]. However, recent advances in deep learning have enabled cost-effective methods, revolutionizing drug design and accelerating the generation of drug-like molecules [38, 1].
Modern drug discovery requires exploring vast chemical spaces under specific constraints to identify molecules that fulfill complex therapeutic objectives. Traditional approaches based on a few scalar properties (e.g., solubility) often fall short in capturing the diversity and complexity of real-world design targets, such as specific structural motifs or functional groups. In contrast, natural language offers a flexible and expressive medium for conveying high-level design intents that encompass functional and structural goals. Consequently, there is increasing interest in the tasks of instruction-based molecule generation and editing [12, 13, 31], which aims to generate novel candidates from natural language instructions or edit existing ones to meet desired instructions, as illustrated in Figure 1.
@@
Figure 1: A generative model works like a drug design specialist, with instruction input and molecule output.
To generate molecules from natural language instructions, prior studies directly use language models on string-based representations like SMILES or SELFIES [48, 27]. This method has been explored by fine-tuning small-scale language models like T5 [12] or adapting large language models [53]. However, such single-modal methods face inherent limitations. Notably, molecular strings lack explicit structural information, such as neighborhood relationships based on edges between atoms [23, 51]. Moreover, representing molecules purely in language can lead to deviation from realistic molecular distributions, as minor string reassembly may produce chemically implausible structures [9]. These issues highlight the need for a multimodal framework for language and graph models, combining robust structural representation with strong instruction-following abilities.
To better represent graph structures in multimodal frameworks, several recent studies have explored powerful graph generation techniques. Among them, graph diffusion methods have shown strong performance and scalability on larger molecules [40, 25]. Recent efforts have introduced guiding the diffusion process using several scalar conditions via predictor-guided [47, 49] or predictor-free approaches [31, 59]. However, when conditioning on natural language, existing methods face notable limitations. While sentences embeddings can replace scalar conditions, they are typically encoded as fixed vectors and injected into node and edge embeddings through simple operations like addition or linear projection. This prevents graph elements from attending to relevant textual tokens, limiting semantic selectivity and hindering fine-grained injection from text, as nodes and edges cannot determine which instruction components are relevant to themselves. Moreover, language representations remain static throughout denoising, preventing interaction and evolving with graph, thereby limiting the model’s ability to refine textual representations and provide guidance tailored to incomplete structures in denoising. These limitations highlight the need for finer-grained and bidirectional integration between language and graph.
Figure 2: Overview of the UTGDiff framework. It incorporates attention bias into the vanilla transformer, forming a unified text-graph transformer that serves as a denoising network for discrete graph diffusion to generate molecular graphs from instructions. Noise decays some nodes and edges into [MASK] during training (forward process), with the reverse process aiming to recover original graphs as the training objective. Sampling starts with a masked graph and iterates $T$ times to reduce noise, where each time it process together with fixed given instruction.
To this end, we propose the Unified Text-Graph Diffusion Model (UTGDiff), a novel conditional graph diffusion framework that unifies language and graph modalities within a single language transformer. By embedding both modalities into a unified transformer, UTGDiff enables more effective token-level text injection, as graph tokens can selectively attend to text tokens via attention. In addition, the instruction representation is continuously updated through attention with graph tokens, allowing it to evolve with molecular structures and contribute back to graph generation. UTGDiff extends masked language models to simulate graph diffusion through three key novel modifications, treating [MASK] tokens as noise in graph, extending the vocabulary with atom-specific tokens, and introducing a loss function to preserve text comprehension for alignment. We further incorporate attention bias into the transformer to model edges.
@@
We conduct extensive experiments on the pre-trained UTGDiff model across instruction-based molecule generation and editing tasks, covering conditions on both molecular structures and properties. Our results demonstrate that, despite its small parameter size, UTGDiff consistently outperforms both language model and graph diffusion baselines under comparable pretraining resources available to us, achieving higher similarity while maintaining a relatively high chemical validity. In summary, our main contributions are:
  • We propose UTGDiff, a novel framework that unifies language and graph modalities within a single language transformer, with graph selectively attending to text and text being refined by graph.
  • We extend language models to simulate graph diffusion by treating [MASK] as graph noise, extending atom-specific tokens, modifying the loss function, and incorporating attention bias for edges.
  • We conduct extensive experiments showing that UTGDiff outperforms language model and graph diffusion baselines in key areas of drug design, achieving higher similarity and relatively high validity under comparable pretraining resources available to us.
2 Related Works
@@
Early works on molecule generation introduced sequence-based models for SMILES [6, 14]. Due to the lack of structural information, recent studies explored graph-based methods, including auto-regressive [16, 55], VAEs [45, 24], GANs [7], and Normalizing Flows [36, 56]. Graph Diffusion is a relatively new approach with superior performance [25, 47], which serves as our foundational framework.
@@
Conditional molecule generation aims to generate a molecule that satisfies a specific property. Early attempts are based on single scalar property such as dipole moment $\mu$ and p-logP scores [30, 21]. More complex conditions involve multi-condition [3, 31] and textual instructions. The latter are addressed by sequence-based generation [12, 5, 28, 41] and multimodal generation employing a graph decoder [46, 32].
@@
In addition to the above instruction-based cross-modal graph generation models, some studies represent molecules and text in cross-modal models for property prediction or text generation [35, 58, 33], employing contrastive alignment, cross-modal projector, or unified text-graph backbone.
@@
Previous Gaussian noise in continuous diffusion models has been shown to be insufficient for discrete data structure. As a solution, researchers introduced the discrete diffusion method, initially discussed in text diffusion [2, 18] and later adapted for graph data [47, 26].
3 Method
3.1 Problem Formulation
The objective of our task is to generate a molecular graph $G=(V,E)$ from a given textual description $S$ specifying molecular functions and properties. The graph consists of $m$ nodes $V\in\mathbb{R}^{m\times a}$ and edges $E\in\mathbb{R}^{m\times m\times b}$ representing node and edge attributes, respectively. Each attribute $v_{i}\in\mathbb{R}^{a}$ and $e_{i,j}\in\mathbb{R}^{b}$ is a one-hot vector corresponding to the categories of the respective nodes and edges. The instruction text is defined as $S=[s_{1},\dots,s_{n}]$, where $s_{i}$ denotes the $i$-th token and $n$ is its length. The goal is to perform conditional graph generation, i.e., sampling $V,E\sim P(V,E\mid S)$.
3.2 Overview of Diffusion Framework
Building upon the above problem setting, We propose a text-conditioned diffusion framework for molecular graph generation, illustrated in Figure 2. As a generative process, the diffusion model iteratively refines random noise into structured data through a sequence of transformations. It involves a forward process which progressively corrupts the data by adding noise according to a noise model, and a reverse process which progressively denoises the data to reconstruct the original molecular structure using a denoising network. The core component of our framework is the unified text-graph denoising transformer, started from a language-pretrained transformer for text comprehension, and enhanced with a [MASK] noise module and graph-specific modeling techniques.
3.3 Forward Process and Noise Model
We adopt the forward diffusion process following Ho et al. [19], where the clean graph $G^{0}$ is progressively corrupted by a noise model $q$, producing a sequence of increasingly noisy graphs $(G^{1},\dots,G^{T})$. It is defined as $q(G^{1:T}\mid G^{0})=\prod^{T}_{t=1}q(G^{t}\mid G^{t-1})$.
Following the discrete diffusion setting in [47], a sequence of transition matrices $(Q^{1},\dots,Q^{T})$ is defined independently for nodes $V$ and edges $E$ as noise model, denoted as $Q_{V}^{t}$ and $Q_{E}^{t}$. Each step in forward process can be expressed as:
@@ 3.3 Forward Process and Noise Model
Under the cumulative transition matrices formulation $\bar{Q}_{V}^{t}=Q_{V}^{1}\dots Q_{V}^{T}$ and $\bar{Q}_{E}^{t}=Q_{E}^{1}\dots Q_{E}^{T}$, the forward process can also be expressed as $q(G^{t}\mid G^{0})=(V\bar{Q}_{V}^{t},E\bar{Q}_{E}^{t})$. With specific transition matrices, the 1-step and t-step marginals are in the closed form, and the posterior at time step t1t-1 admits a closed-form expression:
$\begin{split}q(v_{i}^{t-1}|v_{i},v_{i}^{t})=(v_{i}^{t-1}{Q_{V}^{t}}^{\top}\circ v_{i}\bar{Q}_{V}^{t-1})/(v_{i}\bar{Q}^{t}_{V}{v_{i}^{t}}^{\top})\\ q(e_{ij}^{t-1}|e_{ij},e_{ij}^{t})=(e_{ij}^{t-1}{Q_{V}^{t}}^{\top}\circ e_{ij}\bar{Q}_{V}^{t-1})/(e_{ij}\bar{Q}^{t}_{V}{e_{ij}^{t}}^{\top})\end{split}$ (2)
Graph Diffusion with [MASK] absorbing states
Then we should focus on the formulation of noise model. Traditional methods consider the noise model using the marginal distributions over atom and bond types [47]. While they claimed this makes training easier theoretically and experimentally, such assumptions are tailored to single-modal graph denoising model trained from scratch. In contrast, our goal is to integrate language understanding and graph denoising within a unified language-pretrained transformer, rendering this noise modeling assumption incompatible with our architecture.
@@
Therefore, we employ [MASK] absorbing states originally proposed in text diffusion [2] into our graph diffusion framework. Under this formulation, each element $v_{i}^{t}$ and $e_{i,j}^{t}$ independently decays into a [MASK] token as an absorbing state according to a noise schedule. Following the schedule $\beta(t)=(T-t+1)^{-1}$, the transition matrix $Q^{t}$ is formally expressed as follows for both nodes and edges simultaneously, where $z$ denotes the [MASK] index, and $x,y$ denote categories at consecutive timesteps:
$[Q^{t}]_{x,y}$ $=\left\{\begin{array}[]{l l l}{{1}}&{{\text{if }}}{{x=y=z}}\\ {{1-\beta(t)}}&{{\text{if }}}{{x=y\not=z}}\\ {{\beta(t)}}&{{\text{if }}}{{y=z,x\not=z}}\\ {{0}}&{{\text{otherwise}}}\end{array}\right$. (7)
The feasibility of this method is supported by the interpretation of BERT as a one-step discrete diffusion model [2], where the network is trained to predict the original tokens from [MASK] inputs. As we adopt a language-pretrained transformer familiar with this objective, the formulation aligns with its pretraining goal, thereby facilitating faster convergence of graph diffusion during fine-tuning. It also promotes better integration of text and graph modalities, and helps retain language understanding through direct parameter reuse. Empirical results further validate its effectiveness.
3.4 Reverse Process and Denoising Network
@@
With the initial noisy graph $G^{T}$, the reverse process generates $G^{0}$ iteratively in reversed steps t=T,T1,,0t=T,T-1,...,0, formally represented as $p_{\theta}(G^{0:T-1}\mid G^{T},S)=\prod^{T}_{t=1}p_{\theta}(G^{t-1}\mid G^{t},S)$. It relies on a denoising neural network $\phi_{\theta}$ parameterized by $\theta$, which is trained to invert the forward process by predicting $G^{0}$ from $G^{t}$ conditioned on the instruction $S$ for each $t$, as the product over nodes and edges: $p_{\theta}(G^{0}\mid G^{t},S)=\prod_{v\in V}p_{\theta}(v\mid G^{t},S)\prod_{e\in E}p_{\theta}(e\mid G^{t},S)$.
Then, to calculate the reverse distribution $p_{\theta}(G^{t-1}\mid G^{t},S)$, we adopt $x_{0}$-parameterization with noise back to the clean graph prediction, represented as $p_{\theta}(G^{t-1}\mid G^{t},S)=q(G^{t-1}\mid G^{t},\phi_{\theta}(G^{t},S))$, and marginalized over predictions of node and edge types $\mathcal{V^{\prime}},\mathcal{E^{\prime}}$:
$\begin{split}&\!p_{\theta}(v^{t-1}_{i}\!\mid\!G^{t}\!,\!S)\!=\!\sum_{v\in\mathcal{V^{\prime}}}q(v^{t-1}_{i}\!\mid\!v_{i}\!=\!v,\!G^{t})\;p_{\theta}(v_{i}=v|G^{t},S)\\ &\!p_{\theta}(e^{t-1}_{i,j}\!\mid\!G^{t}\!,\!S)\!=\!\sum_{e\in\mathcal{E^{\prime}}}q(e^{t-1}_{i,j}\!\mid\!e_{i,j}\!=\!e,\!G^{t})\;p_{\theta}(e_{i,j}=e|G^{t},S)\end{split}$ (8)
To enable graph tokens to selectively attend to text tokens, and allow textual representations to be refined by graph structure, we propose the Unified Text-Graph Transformer as our denoising network, which unifies two modalities within a single language-pretrained transformer. Below, we describe two key components of the model.
Molecule Tokenization via Shared Vocabulary
To adapt the language transformer for graph data, we first address the representation of molecular nodes. The key novel design is to treat molecular graph nodes as discrete tokens drawn from a shared vocabulary space, enabling seamless integration with transformer-based language models.
Specifically, we assign a unique textual token (e.g., [C], [O-]) to each atomic category and embed them in the same token space. The initial embeddings are defined as $h^{0}=\text{Emb}(S+V)=[h_{1},\dots,h_{n},h_{n+1},\dots,h_{n+m}]\in\mathbb{R}^{(n+m)\times d_{h}}$, corresponding to $n$ instruction tokens and $m$ graph nodes. The graph node embeddings $h_{n+1},\dots,h_{n+m}$ are derived from the atom tokens (e.g., [[S], [O], [O]] for sulfur oxide).
This tokenization strategy allows graph nodes to be processed identically to language tokens within a unified transformer, enabling attention-based cross-modal integration and consistent parameter sharing. It also prevents compositional ambiguities, such as mistakenly interpreting cobalt (Co) as a combination of carbon and oxygen.
@@
Text Auto-regressive (w.o. pretrain) T5-base 248M 0.731 0.605 0.545 2.48 0.069 0.660
@@
Text Auto-regressive (pretrain) MolT5-base 248M 0.721 0.588 0.529 2.18 0.081 0.772
@@
LLM Chatgpt3.5 (0-shot) - 0.703 0.568 0.517 2.49 0.079 0.721
Chatgpt3.5 (10-shot) - 0.847 0.708 0.624 0.57 0.139 0.887
gpt4 (0 shot) - 0.787 0.577 0.492 1.33 0.055 0.883
gpt4 (10 shot) - 0.872 0.736 0.659 0.66 0.092 0.891
LlaSMol-Mistral 7B 0.853 0.726 0.650 0.70 0.132 0.935